You have $10,000 to invest in a portfolio containing Stock R, Stock S, and a risk-free asset. You must invest all of your money. Your goal is to create a portfolio that has an expected return of 15% and that has only 120% of the risk of the overall market. If Stock R has an expected return of 25% and a beta of 1.6, Stock S has an expected return of 17.5% and a beta of 1.3, and the risk-free rate is 6%, how much money will you invest in Stock R? Explain your answer.

Please show step by step calculations and formulas
Answer #1

The weighted return of stocks in a portfolio is the expected return of the portfolio.

As the beta of Risk-free asset is 0, it is not considered for the below portfolio.

Also, if W1 + W2=1 ,then

W1 = 1-W2

The formula is :

Expected Return = W1R1 + W2R2, where

W1 is the weight of Stock R

R1 Expected return of Stock R

W2 is the weight of Stock S

R2 Expected return of Stock S

15% = 25%(1-W2) + 17.5%(W2)

W2 = 1.333 or Weight of stock S = 1.333

if W1 + W2=1, then W1 is -0.333, or weight of Stock R = -0.333

The weight of Stock R is -0.333 hence it should be shorted to the extent of 33% and that proceeds should be invested in Stock S in the proportion of 1.33, .33 being proceeds from Short selling of Stock R.